# Generative Predictive Control: Flow Matching Policies for Dynamic and Difficult-to-Demonstrate Tasks

### v1

@@ Abstract
Generative control policies have recently unlocked major progress in robotics. These methods produce action sequences via diffusion or flow matching, with training data provided by demonstrations. But despite enjoying considerable success on difficult manipulation problems, generative policies come with two key limitations. First, behavior cloning requires expert demonstrations, which can be time-consuming and expensive to obtain. Second, existing methods are limited to relatively slow, quasi-static tasks. In this paper, we leverage a tight connection between sampling-based predictive control and generative modeling to address each of these issues. In particular, we introduce generative predictive control, a supervised learning framework for tasks with fast dynamics that are easy to simulate but difficult to demonstrate. We then show how trained flow-matching policies can be warm-started at run-time, maintaining temporal consistency and enabling fast feedback rates. We believe that generative predictive control offers a complementary approach to existing behavior cloning methods, and hope that it paves the way toward generalist policies that extend beyond quasi-static demonstration-oriented tasks.
I Introduction and Related Work
Diffusion and flow matching policies have enabled tremendous success in behavior cloning for quasi-static dexterous manipulation tasks [1, 2, 3, 4]. Can generative policies also control systems with fast nonlinear dynamics at high control frequencies, where demonstrations are difficult to come by? We answer this question in the affirmative by introducing Generative Predictive Control (GPC), a supervised learning framework for dynamic and difficult to demonstrate tasks.
The basic idea behind GPC is summarized in Fig. 1. GPC alternates between data collection via Sampling-based Predictive Control (SPC) and policy training via flow matching. The flow matching model bootstraps SPC, allowing for continual performance improvements while maintaining a supervised (e.g., regression) training objective.
I-A Generative Policies for Behavior Cloning
Generative models like diffusion [1] and flow matching [2] have recently gained prominence as powerful policy representations for robotics tasks. These models typically focus on behavior cloning [3, 4], where expert demonstrations serve as training data.
These models offer a few key advantages over other policy representations. Diffusion and flow-matching maintain a high degree of multi-modal expressiveness, allowing for multiple “paths” to achieve the same task [1]. They are also able to consider high-dimensional image observations, which are ubiquitous in modern robotics [5, 6, 7].
Importantly, generative models for behavior cloning are trained in a supervised manner, with clearly defined regression targets. This improves training stability over unsupervised reinforcement learning methods, which can be sensitive to reward specifications, implementation details, and even the random seed used in training [8, 9].
The success of generative behavior cloning, the expressiveness of flow matching and diffusion, and the rise of large foundation models in natural language processing and computer vision have inspired a widespread push toward generalist policies that combine data from multiple tasks [2, 10]. This includes widespread data collection efforts [11] as well as creative tele-operation alternatives [12].
Obtaining demonstration data is a key challenge, and potentially a critical bottleneck in this effort. While creative ways to obtain demonstration data is an area of much active research [12], it is unlikely that demonstrations alone will produce the internet-scale data used to train large vision-language models any time soon. Additionally, some tasks are simply difficult to demonstrate, particularly for robots with fast nonlinear dynamics or unique morphologies.
Figure 1: Generative predictive control is a supervised learning framework for dynamic tasks that are difficult to demonstrate but easy to simulate. First, we generate training data with sampling-based predictive control [13, 14], leveraging advances in massively parallel GPU simulation [15, 16, 17]. We then use this data to train a flow matching policy, which is structured similarly to the behavior cloning policies that have enjoyed tremendous success on quasi-static manipulation tasks [1, 2]. The flow matching policy can then be used to augment the sampling-based controller, further improving performance.
@@ I-A Generative Policies for Behavior Cloning
I-B Sampling-based Predictive Control
At the same time, a very different trend has been gaining traction in the nonlinear optimal control community. Sampling-based Predictive Control (SPC) is an alternative to gradient-based Model Predictive Control (MPC), where a simple sampling procedure is used in place of a sophisticated nonlinear optimizer [13].
@@ I-B Sampling-based Predictive Control
Prominent algorithms in this family include Model Predictive Path Integral Control (MPPI) [14], Predictive Sampling (PS) [18], and Cross-Entropy Methods (CEM) [19]. SPC methods are exceedingly easy to implement, and have been studied for a long time. But recent advances in computing speed and parallel simulation [17, 15, 16] have enabled them to scale to complex problems like in-hand dexterous manipulation [13], legged locomotion [20] and more [21, 18]. On certain tasks, SPC methods offer performance competitive with state-of-the-art reinforcement learning algorithms, even without any offline policy training [13].
SPC is particularly appealing in the context of contact rich robotics tasks, where the stiff dynamics of contact present significant challenges to general-purpose non-convex optimizers [22, 23, 24, 25]. Even the simplest SPC algorithms can outperform specialized contact-implicit optimizers [18], thanks to the fact that they do not require expensive gradients through contact, and the fact that randomization smoothes out contact-related discontinuities [26, 27].
@@
I-C Contribution
In this paper, we highlight a surprisingly deep connection generative policies and SPC algorithms. This connection was first identified by [28] in the context of MPPI: here we extend this connection to a general class of SPC algorithms.
We then leverage this connection to develop a supervised learning framework for dynamic and difficult-to-demonstrate tasks, which we call Generative Predictive Control (GPC). We show how flow-matching policies can be warm-started to encourage temporal consistency and smooth action sequences, and demonstrate how these warm-starts are critical for effective performance on systems with fast dynamics.
We provide an open-source GPC implementation in JAX/MJX [15], which we use to push the scalability limits of our proposed approach. We train GPC policies on systems ranging from a simple inverted pendulum to a 29 degree-of-freedom (DoF) humanoid robot. Our largest and most difficult task (humanoid standup) exposes the limits of this approach: bootstrapping SPC with a trained policy improves performance, but the GPC policy alone is not sufficient.
With this in mind, we provide an extensive discussion of the limitations of GPC and directions for future work to overcome these limitations in Section VIII. These include value function learning, better exploration strategies, and more efficient use of non-optimal samples.
II Background
II-A Problem Formulation
@@
where $x_{\tau}\in\mathbb{R}^{n}$ represents the system state at time step $\tau$, $u_{\tau}\in\mathbb{R}^{m}$ are control actions, $\ell:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ and $\phi:\mathbb{R}^{n}\to\mathbb{R}$ are running and terminal costs, and $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ captures the system dynamics.
For ease of notation, we denote the action sequence as $U=[u_{0},u_{1},\dots,u_{T}]$ and rewrite (1) in compact form as
@@
where both the costs and dynamics constraints are wrapped into the (highly non-convex) objective $J:\mathbb{R}^{m\times T}\times\mathbb{R}^{n}\to\mathbb{R}$.
We will approach this control problem from a Model Predictive Control (MPC) perspective. Rather than attempting to find an optimal closed-form state-feedback policy $u_{\tau}=\kappa(x_{\tau})$, we seek numerical methods to rapidly (and if needed, approximately) solve for the action sequence $U$ from a given initial condition $x_{init}$. At run time, we continually solve from current state estimate, applying only $u_{0}$ before re-planning as quickly as possible.
II-B Sampling-based Predictive Control
@@
In response to these challenges, Sampling-based Predictive Control (SPC) is gaining prominence as a simple and computationally efficient alternative to gradient-based MPC [14, 30, 13, 20, 31]. Instead of relying on the complex machinery of nonlinear optimization, SPC algorithms perform some variation of the following simple procedure:
  • At step $k$, sample $N$ candidate action sequences
    $U^{(i)}\sim\mathcal{N}(\bar{U}_{k-1},\sigma^{2}),\quad i\in[1,N],$ (3)
    typically from a Gaussian proposal distribution [^1] Some SPC methods use a non-isotropic proposal distribution and update the variance of the proposal distribution along with the mean. We focus on an isotropic Gaussian with fixed variance for simplicity of presentation..
@@
    $J^{(i)}\leftarrow J\left(U^{(i)};x_{k-1}\right).$ (4)
  • Use the costs to update the mean action sequence,
    $\bar{U}_{k}\leftarrow\bar{U}_{k-1}+\frac{\sum_{i=1}^{N}g(J^{(i)})(U^{(i)}-\bar{U}_{k-1})}{\sum_{i=1}^{N}g(J^{(i)})},$ (5)
    using some weighting function $g:\mathbb{R}\to\mathbb{R}^{+}$
@@
The ability to parallelize the rollouts (4) presents a key advantage of SPC over traditional gradient-based MPC. The advent of massively parallel GPU-enabled robotics simulators [15, 16, 17] only further reinforces this advantage. Sampling can also reduce the severity of local minima and smooth out stiff dynamics associated with contact [26, 27].
The choice of weighting function $g$ is the key factor distinguishing different SPC algorithms. For instance, MPPI [14] uses a Bolzmann-like exponentially weighted average,
$g_{MPPI}(J)=\exp\left(-J/\lambda\right),$ (6)
where $\lambda>0$ is the temperature parameter. Smaller values of $\lambda$ give more weight to the lowest-cost samples.
In the low-temperature limit we recover predictive sampling [18],
$g_{PS}(J)=\lim_{\lambda\to 0}\exp\left(-J/\lambda\right),$ (7)
Another popular option is CEM [19, 13], which weights the top performing samples equally,
Tsallis-MPPI [32] uses a generalized exponential to provide a middle ground between MPPI and CEM,
$g_{T}(J)=\max(1-(r-1)J/\lambda,0)^{\frac{1}{r-1}},$ (9)
where the original MPPI update is recovered as $r\to 1$.
Remark 1.
SPC performance can benefit significantly from dimensionality reduction via spline representations, e.g., representing $U=[u_{0},u_{1},\dots,u_{T}]$ with $T^{\prime}\ll T$ knot points. While our implementation focuses on only the simplest zero-order-hold interpolation, the GPC framework is compatible with more sophisticated schemes such as Fourier features or higher-order splines [18].
II-C Generative Modeling
II-C1 Flow Matching
Flow matching generative models [7] focus on a seemingly very different problem: produce a sample $\mathbf{x}$ from a probability distribution $p(\mathbf{x})$. In the typical problem setting, we do not have access to $p(\mathbf{x})$ in closed form, but we do have samples from this data distribution.
The basic idea is to consider a probability density path $p_{t}(\mathbf{x})$. This path flows from an easy-to-sample proposal distribution $p_{0}(\mathbf{x})=\mathcal{N}(0,I)$ at $t=0$ to the data distribution at $t=1$. Flow matching generative models learn a time-varying vector field $v_{\theta}(\mathbf{x},t)$ that moves samples from $p_{0}(\mathbf{x})$ to $p_{1}(\mathbf{x})$, where $\theta$ denote learnable parameters.
The flow matching network $v_{\theta}$ is trained via standard (stochastic) gradient descent methods on
@@ II-C1 Flow Matching
where
$\mathcal{L}_{FM}\left(\theta;\mathbf{x}_{0},\mathbf{x}_{1},t\right)=\\ \left\|v_{\theta}\big{(}t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0},t\big{)}-(\mathbf{x}_{1}-\mathbf{x}_{0})\right\|^{2}$ (11)
is the flow matching loss. The expectation in (10) taken over $t\sim\mathcal{U}(0,1)$, $\mathbf{x}_{0}\sim p_{0}(\mathbf{x})$, and $\mathbf{x}_{1}\sim p_{1}(\mathbf{x})$. Each of these are easy to acquire, since we already have data points $\mathbf{x}_{1}$ and $p_{0}(\mathbf{x})=\mathcal{N}(0,I)$ is easy to sample from. Intuitively, $\mathcal{L}_{FM}$ attempts to push samples in straight line from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$.
At inference time, we generate new samples by starting with $\mathbf{x}\sim p_{0}(\mathbf{x})$ and simulating the Ordinary Differential Equation (ODE)
$\dot{\mathbf{x}}=v_{\theta}(\mathbf{x},t).$ (12)
Any off-the-shelf ODE solver can be used for this purpose, but in many cases a simple forward Euler scheme
$\mathbf{x}_{t+\delta t}=\mathbf{x}_{t}+\delta tv_{\theta}(\mathbf{x}_{t},t),$ (13)
is sufficiently performant.
II-C2 Diffusion
Flow matching is equivalent (under some technical conditions) to diffusion-based generative modeling [33]. Diffusion models also consider a series of probability distributions flowing from an initial Gaussian to the data distribution. But rather than being parameterized by a time $t$, these are typically parameterized by additive noise $\sigma$,
@@ II-C2 Diffusion
For large $\sigma$, $p_{\sigma}(\mathbf{x})$ is dominated by noise and approaches an easy-to-sample Gaussian. For small $\sigma$, $p_{\sigma}(\mathbf{x})$ approaches the data distribution.
Diffusion models estimate the score
$s_{\theta}(\mathbf{x},\sigma)\approx\nabla_{\mathbf{x}}\log p_{\sigma}(\mathbf{x})$ (15)
at various noise levels by learning to remove noise added to the original data [6, 5]. Once we have a trained score network, we can use it sample from $p_{\sigma}$ using Langevin dynamics
@@
III SPC is Online Generative Modeling
In this section, we establish a formal connection between SPC and generative modeling. Specifically, we show that the SPC update (5) is a monte-carlo estimate of the score of a noised target distribution. This connection was first identified for the case of MPPI in [28] and used to develop Dial-MPC, a multi-stage SPC algorithm for legged locomotion, in [20]. Here we extend this connection to generic SPC algorithms with updates of the form (5).
@@ III SPC is Online Generative Modeling
which is determined by the algorithm-specific weighting function $g(\cdot)$ introduced in Sec. II-B above. In the spirit of score-based diffusion (Sec. II-C2) we define the noised target distribution

### v2

# Generative Predictive Control: Flow Matching Policies for Dynamic, Difficult-to-Demonstrate Tasks

@@ Abstract
Generative control policies have recently unlocked major progress in robotics. These methods produce action sequences via diffusion or flow matching, with training data provided by demonstrations. But existing methods come with two key limitations: they require expert demonstrations, which can be difficult to obtain, and they are limited to relatively slow, quasi-static tasks. In this paper, we leverage a tight connection between sampling-based predictive control and generative modeling to address each of these issues. In particular, we introduce generative predictive control, a supervised learning framework for tasks with fast dynamics that are easy to simulate but difficult to demonstrate. We then show how trained flow-matching policies can be warm-started at inference time, maintaining temporal consistency and enabling high-frequency feedback. We believe that generative predictive control offers a complementary approach to existing behavior cloning methods, and hope that it paves the way toward generalist policies that extend beyond quasi-static demonstration-oriented tasks.
@@
Keywords: Generative models, flow matching, sampling-based predictive control
1 Introduction and Related Work
Diffusion and flow matching policies have enabled tremendous success in behavior cloning for quasi-static manipulation [1, 2, 3, 4]. Can generative policies also control systems with fast nonlinear dynamics at high control frequencies, where demonstrations are difficult to come by? In this paper, we answer this question in the affirmative by introducing generative predictive control (GPC), a supervised learning framework for dynamic and difficult-to-demonstrate tasks.
Figure 1: Generative predictive control is a supervised learning framework for dynamic tasks that are difficult to demonstrate but easy to simulate. First, we generate training data with sampling-based predictive control [5, 6, 7], leveraging advances in massively parallel GPU simulation [8, 9, 10]. We then use this data to train a flow matching policy, which in turn provides additional high-quality samples. This results in better training data for subsequent iterations, in a virtuous cycle.
Fig. 1 summarizes our approach. GPC alternates between data collection via sampling-based predictive control (SPC) and policy training via flow matching. The flow model provides extra samples for SPC, enabling continual performance improvement while maintaining a supervised (regression) objective.
Generative Policies: Diffusion [1] and flow matching [2] have recently gained prominence as powerful policy representations for robotics. These models typically focus on behavior cloning [3, 4], where expert demonstrations serve as training data. Generative policies have the key advantage of multi-modal expressiveness, allowing multiple “paths” to the same goal [1]. They also provide a natural choice for handling image data [11, 12, 13].
Importantly, generative policies are trained in a supervised manner, with clearly defined regression targets. This improves training stability over unsupervised reinforcement learning [14, 15], but requires demonstraions. Obtaining sufficient demonstration data is a key challenge, particularly for large generalist policies [2, 16, 17]. While creative ways to obtain this data are an area of active research [18], it is unlikely that demonstrations alone will produce the internet-scale data used to train large vision-language models any time soon. Furthermore, some tasks are simply difficult to demonstrate, particularly for robots with fast nonlinear dynamics or unique morphologies.
@@ 1 Introduction and Related Work
Sampling-based Predictive Control (SPC): At the same time, a very different trend has been gaining traction in the nonlinear optimal control community. SPC is an alternative to gradient-based model predictive control (MPC), where a simple sampling procedure is used in place of a sophisticated nonlinear optimizer [19, 20, 21]. Algorithms in this family include model predictive path integral control (MPPI) [6], predictive sampling (PS) [7], and cross-entropy methods (CEM) [22]. SPC algorithms are exceedingly easy to implement, and have been studied for a long time. But recent advances in computing speed and parallel simulation [8, 9, 10] have enabled them to scale to complex problems like in-hand dexterous manipulation [5], legged locomotion [23] and more [24, 7].
@@
Our Contributions: In this paper, we highlight a surprisingly deep connection between generative policies and SPC. This connection was first identified by [25] in the context of MPPI: here we extend this connection to a general class of SPC algorithms.
We then leverage this connection to propose a supervised learning framework for difficult-to-demonstrate tasks, GPC. We show how flow-matching policies can be warm-started to encourage temporal consistency, and demonstrate that these warm-starts are critical for high-frequency feedback. To push the scalability limits of GPC, we train policies on systems ranging from a simple inverted pendulum to a humanoid robot, and evaluate effectiveness in simulation. Our largest and most difficult task (humanoid standup) exposes these scalability limits: seeding SPC with a GPC policy is effective, but applying the policy directly is not. With this in mind, we provide an extensive discussion of limitations and directions for future work.
2 Background
@@
where $x_{\tau}\in\mathbb{R}^{n}$ represents the system state at time step $\tau$, $u_{\tau}\in\mathbb{R}^{m}$ are control actions, $\ell(\cdot,\cdot)$ and $\phi(\cdot)$ are running and terminal costs, and $f(\cdot,\cdot)$ captures the system dynamics. For simplicity, we denote the $T$-length action sequence as $U=[u_{0},u_{1},\dots,u_{T}]$ and rewrite (1) in compact form as
@@
where both the costs and dynamics constraints are wrapped into the (highly non-convex) objective.
2.1 Sampling-based Predictive Control
@@
In response to these challenges, SPC is gaining prominence as a simple and computationally efficient alternative to gradient-based MPC [5, 6, 23, 27, 28]. Instead of relying on the complex machinery of nonlinear optimization, SPC algorithms perform some variation on the following simple procedure:
  1. 1.
    At step $k$, sample $N$ candidate action sequences from a Gaussian proposal distribution [^1] Some SPC methods use a non-isotropic proposal distribution and update the variance of the proposal distribution along with the mean. We focus on an isotropic Gaussian with fixed variance for simplicity..
    $U^{(i)}\sim\mathcal{N}(\bar{U}_{k-1},\sigma^{2}),\quad i\in[1,N].$ (3)
  2. 2.
@@
    $J^{(i)}=J\left(U^{(i)};x_{k-1}\right).$ (4)
  3. 3.
    Update the mean action sequence according to some weighting function $g:\mathbb{R}\to\mathbb{R}^{+}$,
    $\bar{U}_{k}=\bar{U}_{k-1}+\frac{\sum_{i=1}^{N}g(J^{(i)})(U^{(i)}-\bar{U}_{k-1})}{\sum_{i=1}^{N}g(J^{(i)})}.$ (5)
  4. 4.
@@
The weighting function $g$ distinguishes different SPC algorithms. As detailed in Appendix A, different choices give rise to MPPI [6], predictive sampling [7], CEM [22, 5], and more [29].
The ability to parallelize rollouts (4) is a key advantage over gradient-based MPC. The advent of massively parallel GPU simulators [8, 9, 10] further reinforces this advantage. Sampling can also reduce the severity of local minima and smooth out stiff dynamics associated with contact [30, 31].
2.2 Generative Modeling
Generative modeling considers a seemingly very different problem: produce a sample $\mathbf{x}$ from a probability distribution $p(\mathbf{x})$. In the typical setting, we do not have access to $p(\mathbf{x})$ in closed form, but we do have samples (training data) from this data distribution.
Flow Matching: Flow matching [13] is based on the idea of a probability density path $p_{t}(\mathbf{x})$. This path flows from an easy-to-sample distribution $p_{0}(\mathbf{x})=\mathcal{N}(0,I)$ at $t=0$ to the data distribution at $t=1$. Flow matching methods learn a vector field $v_{\theta}(\mathbf{x},t)$ that moves samples along this path.
The flow network $v_{\theta}$ is trained via standard (stochastic) gradient descent methods on
@@ 2.2 Generative Modeling
where $\theta$ are learnable parameters (e.g., network weights and biases) and
$\mathcal{L}_{FM}\left(\theta;\mathbf{x}_{0},\mathbf{x}_{1},t\right)=\left\|v_{\theta}\big{(}t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0},t\big{)}-(\mathbf{x}_{1}-\mathbf{x}_{0})\right\|^{2}.$ (7)
The expectation in (6) is taken over $t\sim\mathcal{U}(0,1)$, $\mathbf{x}_{0}\sim p_{0}(\mathbf{x})$, and $\mathbf{x}_{1}\sim p_{1}(\mathbf{x})$. Each of these are easy to sample, since we already have data points $\mathbf{x}_{1}$, and sampling $p_{0}(\mathbf{x})$ is trivial. Intuitively, $\mathcal{L}_{FM}$ pushes samples in a straight line from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$. At inference time, we first sample $\mathbf{x}\sim p_{0}(\mathbf{x})$, then integrate $\dot{\mathbf{x}}=v_{\theta}(\mathbf{x},t)$ from $t=0$ to $t=1$, typically with a simple explicit Euler scheme.
Diffusion: Flow matching is equivalent (under some technical conditions [32]) to diffusion-based generative modeling [11, 12]. Diffusion models also consider a series of probability distributions flowing from an initial Gaussian to the data distribution. But rather than being parameterized by a time $t$, these are typically parameterized by additive noise $\sigma$,
@@
For large $\sigma$, $p_{\sigma}(\mathbf{x})$ approaches an easy-to-sample Gaussian. For small $\sigma$, $p_{\sigma}(\mathbf{x})$ approaches $p(\mathbf{x})$.
Diffusion models learn the score $s_{\theta}(\mathbf{x},\sigma)\approx\nabla_{\mathbf{x}}\log p_{\sigma}(\mathbf{x})$ by removing noise added to the original data [11, 12]. We can then use $s_{\theta}$ to sample from $p_{\sigma}$ using Langevin dynamics
@@
3 SPC is Online Generative Modeling
In this section, we establish a formal connection between SPC and generative modeling. Specifically, we show that the SPC update (5) is a Monte Carlo estimate of the score of a noised target distribution. This connection was first identified for the case of MPPI in [25] and used to develop Dial-MPC, a multi-stage SPC algorithm for legged locomotion [23]. Here we extend this connection to generic SPC algorithms with updates of the form (5).
@@ 3 SPC is Online Generative Modeling
which is determined by the algorithm-specific weighting function $g(\cdot)$ introduced in Sec. 2.1 above. In the spirit of score-based diffusion [12] we define the noised target distribution